?(a) Find the? t-value such that the area in the right tail is 0.20 with 21 degrees of freedom. ?(Round to three decimal places as? needed.)

?(b) Find the? t-value such that the area in the right tail is 0.15 with 16 degrees of freedom.?(Round to three decimal places as? needed.)

?(c) Find the? t-value such that the area left of the? t-value is 0.005 with 19 degrees of freedom.? [Hint: Use? symmetry.] ?(Round to three decimal places as? needed.)

(d) Find the critical? t-value that corresponds to 90% confidence. Assume 21 degrees of freedom ?(Round to three decimal places as? needed.)
Answer #1

a_)

from excel: t value =t.inv(1-0.20,21)=0.859

b)t value =T.INV(1-0.15,16) =1.071

c) t value =t.inv(0.005,19)=-2.861

d)

t value =t.inv((1+0.9)/2,21) =1.721